OpenStudy (anonymous):

How would I do this problem? 3^5/3

OpenStudy (anonymous):

3 to the power of 5 divided by 3. so first what you want to do is your exponet's 3^5 power which is 243 and then your division. so whats 243/3 ?

OpenStudy (anonymous):

I believe I have to simplify

OpenStudy (k_lynn):

\[3^{5}\div3\]

OpenStudy (k_lynn):

You basically plug into your calculator, 3x3x3x3x3 or \[3^{5}\] then you divide the product by 3.

OpenStudy (anonymous):

Well said. ^

OpenStudy (k_lynn):

^.^ Thanks.

OpenStudy (diamondboy):

\[3^{5}/3\]

OpenStudy (diamondboy):

is dat your question?

OpenStudy (anonymous):

Thanks for the help but what im doing is differently here's a sample problem from my notes; 16^3/4 and then it led to (16^1/4)^3 and that led to (2)^3=8

OpenStudy (anonymous):

I'm entirely lost

OpenStudy (diamondboy):

\[x ^{a/x} = \sqrt[x]{1}^{a}\]

OpenStudy (diamondboy):

that is d method u need to use

OpenStudy (diamondboy):

(3^1/3)^5 \[(\sqrt[3]{3})^{5}\]

OpenStudy (ilovecake):

answer is 81